A 100.0 −mL buffer solution is 0.100 M in NH3 and 0.130 M in NH4Br.

Part A: What mass of HCl can this buffer neutralize before the pH falls below 9.00?

Part B: If the same volume of the buffer were 0.265 M in NH3 and 0.395 M in NH4Br, what mass of HCl could be handled before the pH falls below 9.00?
Answer #1

The pOH of the buffer is calculated as shown below: POH=pK, +108 sati 0.130M = -log(1.8x10)+log 100 M = 4.74 +0.114 = 4.854 T

The pOH is calculated as shown below: (salt) pOH =pK, +log [base] 13+xmmoles 14-9 =4.74 +log F109 10-x mmoles 13+xmmoles 0.26

[salt] pOH =pK, +logi [base] 39.5+xmm oles 14-9=4.74 +log 26.5-x mmoles 39.5+xmmoles =10926.5-x mm oles 39.5+xmmoles 26.5-x m